1. What is the equivalent resistance of two parallel resistors, one of 40 Ω and the other 60 Ω? What is the equivalent capacitance oftwo parallel capacitors, one of 10μF and the other 15 μΕ? 2. a) You pick up a resistor and notice its color bands are BLUE, YELLOW, and What is the value of this resistor? RED. 3. b) and another has BROWN, BLACK, BROWN. What is its resistance?
Answer #1

1)

In parallel,

1/Req = 1/R1 + 1/R2

1/Req = 1/40 + 1/60

1/Req = (3+2)/120

1/Req = 5/120

Req = 120/5

= 24 ohm

Answer: 24 ohm

2)

In Parallel,

Ceq = C1 + C2

= 10 uF + 15 uF

= 25 uF

Answer: 25 uF

I am allowed to answer only 1 question at a time